Science and medicine. A bicyclist rode into the country for 5 h. In returning, her
speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?

Answer by ankor@dixie-net.com(22740)

A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h
faster and the trip took 4 h. What was her speed each way?
:
Let s = her speed into the country (To commune with nature, no doubt)
Then (s+5) = her speed returning from the country (Chased by a bear possibly!)
:
Distance = time * speed
:
Dist out = Dist back
5s = 4(s+5)
5s = 4s + 20
5s - 4s = 20
s = 20 mph out, 25 mph back
:
:
Check using distance:
out: 5(20) = 100 mi
back 4(25) = 100 mi